A contractor mixes concrete from bags of pre-mix for small jobs. How many bags with 11% cement should he mix with 8 bags of 13% of cement to produce a mix containing 12% cement.
I tried setting up a table based of the example in my text book and I still can't get it to work.
Thanks

Answer by scott8148(6628)

because 12 is exactly midway between 11 and 13, the answer is 8

the equation is ___ x(11%) + 8(13%) = (x + 8)(12%) ___ multiply by 100 to clear the percents

11x + 104 = 12x + 96

8 = x